Operations Report – January 10th



Crew 272 Operations Report 10-1-2023

SOL: 9

Name of person filing report: Mason Kuhn

Non-nominal systems: Wiring in EVA Suit 7

Notes on non-nominal systems: Due to the RAM being closed to the crew this morning, EVA, Hab operations, and time-sensitive personal items, I did not have time to repair Suit 7. This will be completed during sol 10.


Spirit rover used: Yes

Hours: 207.1

Beginning charge: 100%

Ending charge: 69%

Currently charging: Yes

Opportunity rover used: No

Hours: 111.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 219.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 253.9

Beginning charge: 100%

Ending charge: 58%

Currently charging: Yes

General notes on rovers: Spirit and Perseverance were used on the EVA. After returning to the Hab, per mission support request, all 4 rover positions were slightly adjusted back to their correct parking spots. A crewmate noticed that the rear-left taillight on Perseverance was non-nominal.

Summary of Hab operations: Around 8:00 pm, despite already being roped shut, the outer front airlock door flew open. The crew was prepared and scrambled to get ready with EVA Suits, boots, and other essential items. The problem was very promptly addressed, and all crew members are safe.

WATER USE: 19.6 gallons

Water (static tank): 206.9 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal (with the exception of a ~15-minute anomaly last night around 8:40 pm)

Summary of suits and radios: Suits 1, 8, 9, and 10 were used during the SAR EVA to Kissing Camel Ridge. All suits performed nominally with airflow and minimal fogging. All suits have been cleaned and are currently charging. There were no issues with radios or headsets on the EVA. All radios that were used have been charged and removed from the charging dock.

Summary of GreenHab operations: Monitored and watered GreenHab crops and experimental crops. Placed new shovels and broom in the GreenHab.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 10g beans, 80g carrots

Summary of Science Dome operations: No crew activity, there was Mission Support activity.

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: The items that were dropped off on Mars were retrieved from the RAM and put into their proper location (GreenHab, Kitchen, EVA Suit Cabinet). The items that were not retrieved and are still in the RAM are the goggles for the Sci Dome and the Observatory Items. After the supply trip was made, the Hab had an emergency situation of a blown open airlock door. This was promptly taken care of and all crew members are safe, but the storm prohibited a crew member from going to back to retrieve the remaining items. They will be taken care of when the weather permits.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Where is the proper location for Observatory items?

Best Regards,

Mason Kuhn

Crew Engineer

Copyright © The Mars Society. All rights reserved. | Main Site